Please help i need to get these right in order to get an A PLEASE PLEASE PLEASE

1. What is the Kc for the following reaction at equilibrium at 1000 ∘C if [CO]= 0.50 M , [H2]= 0.32 M , [CH4]= 1.6 M , and [H2O]= 2.8 M ? CO(g)+3H2(g)⇌CH4(g)+H2O(g)

2. The equilibrium constant, Kc, for the reaction of H2 and N2 is 1.7×102 at 225 ∘C.
3H2(g)+N2(g)←−→2NH3(g)

If the system at equilibrium contains 0.19 M H2 and 0.026 M N2, what is the concentration of NH3?

Express your answer using two significant figures.

3. For each of the following Kc values, indicate whether the equilibrium mixture contains mostly reactants, mostly products, or similar amounts of reactants and products.

Part A

N2(g)+O2(g) ←−→ 2NO(g) Kc=1×10−30

A) The equilibrium mixture contains mostly reactants.
B) The equilibrium mixture contains mostly products.
C) The equilibrium mixture contains similar amounts of reactants and products.

Part B

H2(g)+Br(g) ←−→ 2HBr(g) Kc=2.0×1019

A) The equilibrium mixture contains mostly reactants.
B) The equilibrium mixture contains mostly products.
C) The equilibrium mixture contains similar amounts of reactants and products
Answer #1

1. What is the Kc for the following reaction at equilibrium at 1000 ∘C if [CO]= 0.50 M , [H2]= 0.32 M , [CH4]= 1.6 M , and [H2O]= 2.8 M ? CO(g)+3H2(g)⇌CH4(g)+H2O(g)

Kc = [CH4][H2O] / [CO][H2]^3

Kc = (1.6*2.8)/(0.5*(0.32^3)) = 273.4375

2. The equilibrium constant, Kc, for the reaction of H2 and N2 is 1.7×102 at 225 ∘C.
3H2(g)+N2(g)←−→2NH3(g)

If the system at equilibrium contains 0.19 M H2 and 0.026 M N2, what is the concentration of NH3?

K = [NH3]^2 / [H2]^3 [ N2]

1.7*10^2 = (NH3)^2 * (0.19^3 * 0.026)

NH3 = sqrt(0.0303) = 0.174 M

3)

A)

K is too low, this favour reactant formation

B)

K is too high, this favours products formation